S Halliday Fundamentals of yscs, 10e Cakules based Physics 1&E (P4 201 Chapter 04, Problem 064 partide moves horizentally in unilorm circular motion, over a horlontal ay glane. Al one instant, i meves throwgh the point at coordinutes (3.0m, 4 90 m) with aelocity of s.50 7 m/s and an acceleration of +13.27 ms. what are the Ca) and() crdinates af the center of the droda (a) Number (b) Number Unit 0 ef a used search OEL SPACE 7 9 0 5
Answer #1

We know that,

a = v^2 / r

r = v^2 / a = (5.50)^2 / 13.2 = 2.29 m

Since r is downward,

x = 3.70 m

y = 4.90 - 2.29 = 2.61 m